Convert a chromel-alumel thermocouple measurement of 3.100 millivolts (mV) to degress Celsius. For a chromel-alumel type K thermocouple, 3.100 mV is between the entries of 2.667 (at 150°F) and 3.819 (at 200°F). Assume a linear relation between the two.

1
Expert's answer
2021-11-06T01:37:50-0400

3.819-2.667=1.152mV

200-150=50°F

"\therefore1.152mV=50°F\newline1mV=?"

"=\frac{1*50}{1.152}=43.4028"

3.100mV=3.100*43.4028=134.55°F

Converting °F to °C

"\frac{(F-32)*5}{9}=C"

"\frac{(134.55-32)*5}{9}=56.97"

=56.97°C